OpenStudy (anonymous):

solving linear equations x-2/3=2x+1/4

OpenStudy (anonymous):

multiply both sides by 12

OpenStudy (anonymous):

is it x - (2/3) = 2x + (1/4) or is it (x - 2)/3 = (2x + 1)/4 ???

OpenStudy (anonymous):

(x - 2)/3 = (2x + 1)/4

OpenStudy (anonymous):

(x - 2)/3 = (2x + 1)4 1/3x - 2/3 = 1/2x + 1/4 1/3x - 1/2x = 1/4 + 2/3 (multiply everything by 12) 4x - 6x = 3 + 8 -2x = 11 x = -11/2